Problem 3-12 Predetermined Overhead Rate; Disposing of Underapplied or Overapplied Overhead [LO3- 4] Luzadis Company makes furniture using the latest automated technology. The company uses a job-order costing system and applies manufacturing overhead cost to products on the basis of machine-hours. The predetermined overhead rate was based on a cost formula that estimates $880,000 of total manufacturing overhead for an estimated activity level of 88,000 machine-hours During the year, a large quantity of furniture on the market resulted in cutting back production and a buildup of furniture in the companys warehouse. The companys cost records revealed the following actual cost and operating data for the year: Machine-hours Manufacturing overhead cost Inventories at year-end: 79,000 s 836,000 Raw materials Work in process (includes overhead applied of $71,100) Finished goods (includes overhead applied of $134,300) $ 15, 000 $ 171,900 $ 324,700 $ 1,413,400 Cost of goods sold (includes overhead applied of $584,600) Required: 1. Compute the underapplied or overapplied overhead 2. Assume that the company closes any underapplied or overapplied overhead to Cost of Goods Sold. Prepare the appropriate journal entry 3. Assume that the company allocates any underapplied or overapplied overhead proportionally to Work in Process, Finished Goods, and Cost of Goods Sold. Prepare the appropriate journal entry 4. How much higher or lower will net operating income be if the underapplied or overapplied overhead is allocated to Work in Process, Finished Goods, and Cost of Goods Sold rather than being closed to Cost of Goods Sold? Complete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Compute the underapplied or overapplied overhead Required 1 Required 2Required 1Required 2 Required 3 Required 4 Assume that the company closes any underapplied or overapplied overhead to Cost of Goods Sold. Prepare the appropriate journal entry. (If no entry is required for an transaction/event, select No journal entry required in the first account field. .. View transaction list Journal entry worksheet Record the entry to close the balance in the manufacturing overhead account to the cost of goods sold account. Note: Enter debits before credits. Event General Journal Debit Credit Record entry Clear entry View general journalRequired 1 Required 2Required 3 Required 4 Assume that the company allocates any underapplied or overapplied overhead proportionally to Work in Process, Finished Goods, and Cost of Goods Sold. Prepare the appropriate journal entry. (If no entry is required for a transaction/event, select No journal entry required in the first account field.) View transaction list Journal entry worksheet Record the allocation of the underapplied/overapplied overhead to various accounts. Note: Enter debits before credits. Event General Journal Debit Credit Record entry Clear entry View general journalComplete this question by entering your answers in the tabs below. Required 1Required 2 Required 3Required 4 How much higher or lower will net operating income be if the underapplied or overapplied overhead is allocated to Work in Process, Finished Goods, and Cost of Goods Sold rather than being closed to Cost of Goods Sold? Net operating income will be overhead is allocated rather than closed entirely to cost of goods sold if the
Answer #1

1) Overhead rate = 880000/88000 = $10 per machine hour

Applied overhead = 79000*10 = $790000

Actual overhead = $836000

Under applied overhead = (790000-836000) = $46000

2) Journal entry

Date account and explanation debit credit
Cost of goods sold 46000
Manufacturing overhead 46000
(To record under applied overhead)

3) Journal entry

Date account and explanation debit credit
Work in process (46000*9%) 4140
Finished goods (46000*17%) 7820
Cost of goods sold (46000*74%) 34040
Manufacturing overhead 46000
(To record under applied overhead)

4) Net operating income will higher (4140+7820) = $11960 if the under applied overhead is allocated rather than closed entirely to Cost of goods sold